OpenStudy (laddiusmaximus):

a ball is tossed up vertically and returns to earth 1 s later. what is the initial velocity and how high did it go?

OpenStudy (anonymous):

\[h(t)=-16t^2+v_0t\] and i guess we know that \[h(1)=0\] so we have \[-16+v_0=0\] \[v_0=16\] initial velocity is 16 feet per second ( if we are working in feet)

OpenStudy (rulnick):

Yes, I believe 16 ft/s is correct, and the ball reached a max height of 4 ft.

OpenStudy (laddiusmaximus):

its in meters

OpenStudy (laddiusmaximus):

i cant figure out the equation for this to work

OpenStudy (rulnick):

4.9 m/s initial velocity, 1.22 m max height

OpenStudy (rulnick):

I'll help with the equation, hang on.

OpenStudy (laddiusmaximus):

1.22 isnt right

OpenStudy (rulnick):

The equation would be x(t) = x(0) + v(0)t + (1/2)at^2. a is -9.8 m/s^2. x(0) is assumed to be 0 if we are tossing from the ground. Good so far?

OpenStudy (rulnick):

So x(t) = v(0)t-4.9t^2. We are interested in x(1)=0 at t=1 s, so ...

OpenStudy (rulnick):

0 = v(0) (1) - 4.9 (1)^2 0 = v(0) -4.9 v(0) = 4.9 (m/s).

OpenStudy (rulnick):

The max height is reached at t=1/2: x(1/2) = 4.9(1/2) -4.9(1/2)^2, so ...

OpenStudy (rulnick):

x(1/2) = 4.9 (1/4) = 1.225 (meters).

OpenStudy (laddiusmaximus):

got it. its confusing. why did you use 4.9?

OpenStudy (rulnick):

4.9 = (1/2) (9.8).

OpenStudy (rulnick):

You think it's supposed to be something other than 1.225 meters?

OpenStudy (laddiusmaximus):

no its right.

OpenStudy (laddiusmaximus):

i put 1.22